OpenStudy (xxstudyingstuffxx):

Exactly 2/5 of the university's students live off campus. Of those students, 3/10 live at home. What fraction of the university's students live at home? Reduce your answer to lowest terms. Will fan and medal.

OpenStudy (xxstudyingstuffxx):

I think it is 3/25. Am I right?

OpenStudy (xxstudyingstuffxx):

@AloneS @ILovePuppiesLol

OpenStudy (plainntall):

What fraction of students live off campus? 2/5 Of those students 3/10 of them live at home so that means 3/10 of 2/5 of students live at home or (3./10)(2/5) = 6/50 of all students live at home. If we factor out 2 we have (2*3)/(2*25) = 3/25 so yes you are correct :)

OpenStudy (xxstudyingstuffxx):

Oh. I did it this way: 500 students total 300 are on campus 200 are off campus 60 of the off campus people are at home 60 is 3/25 of 500. Thanks, though!